Given the training set below:

Age

Income

Student

Credit_rating

Buys_computer

<=30

high

no

fair

no

<=30

high

no

excellent

no

31…40

high

no

fair

no

>40

medium

no

fair

yes

<=30

low

no

fair

yes

>40

high

no

fair

no

>40

low

yes

fair

yes

>40

low

yes

excellent

no

31…40

low

yes

excellent

yes

<=30

medium

no

fair

no

<=30

low

yes

fair

yes

>40

medium

yes

fair

yes

<=30

medium

yes

excellent

no

31…40

medium

no

excellent

no

31…40

high

yes

fair

yes

>40

medium

no

excellent

yes

The information gain for attribute Credit_rating is:

Group of answer choices

1. 0.049

2. 0.244

3. 0.261

4. 0.488

5. 0.004
Answer #1

In column Credit_rating, we have 6 excellents and 10 fairs. So, Entropy of Credit_rating is
-\frac{6}{16}log_2(\frac{6}{16})-\frac{10}{16}log_2(\frac{10}{16})
= 0.954

Now, we split Credit_rating based on Buys_computer. When Buys_computer is yes, we have 6 fairs and 2 excellents. When Buys_computer is no, we have 4 fairs and 4 excellents.
Entropy of Buys_computer=yes is
-\frac{6}{8}log_2(\frac{6}{8})-\frac{2}{8}log_2(\frac{2}{8})

= 0.811

Entropy of Buys_computer=no is
-\frac{4}{8}log_2(\frac{4}{8})-\frac{4}{8}log_2(\frac{4}{8})

= 1

Entropy of Credit_rating based on Buys_computer is
0.5*0.811+0.5*1
= 0.9055

So, Information gain from Credit_rating is 0.954 - 0.9055 = 0.0485 ~ 0.049

So, the answer is 0.049